# Enhancing Time Series Forecasting via Multi-Level Text Alignment with LLMs

### v1

@@ 4.1 Long-term Forecasting
ETTh2 0.397 0.431 0.446 0.464 0.366 0.407 0.397 0.421 0.605 0.538 0.415 0.431
ETTm2 0.262 0.324 0.292 0.343 0.276 0.324 0.293 0.335 0.316 0.368 0.296 0.343
Weather 0.207 0.263 0.359 0.275 0.235 0.270 0.238 0.275 0.241 0.283 0.242 0.279
ECL 0.190 0.288 0.182 0.277 0.172 0.264 0.176 0.269 0.180 0.280 0.180 0.273
Traffic 0.409 0.310 0.438 0.312 0.432 0.303 0.440 0.310 0.447 0.313 0.430 0.305
Table 3: Zero-shot learning results on ETT datasets. Lower values indicate better performance. bold: the best, underline: second best.
@@
4.3 Zero-shot forecasting
Setups. Beyond few-shot learning, LLMs also show promise as effective zero-shot learners. In this section, we evaluate the zero-shot learning capabilities of the multi-level text-aligned LLM. Specifically, we assess how well the model performs on one dataset after being optimized on another. Similar to the few-shot learning setup, we use the long-term forecasting protocol and evaluate various cross-domain scenarios utilizing the ETT datasets.
Results. The brief results are presented in Table 3, with full results in Table 6. Our model demonstrates performance that is comparable to or surpasses other baselines. In data-scarce scenarios, our model significantly outperforms other LLM-based models, consistently providing better forecasts. Both our model and TimeLLM [10] outperform traditional baselines, likely due to cross-modality alignment, which more effectively activates LLMs’ knowledge transfer and reasoning capabilities for time series tasks. Additionally, our multi-level aligned embeddings better align language cues with temporal components of time series, enabling superior zero-shot forecasting performance compared to TimeLLM.
4.4 Model analysis
Table 4: Comparison of different variants for long-term and few-shot forecasting.
Variant Long-term Forecasting Few-shot Forecasting
ETTm1-96 ETThm1-192 ETTm1-96 ETTm1-192
@@ 4.4 Model analysis
B.1 only trend alignment 0.184 0.283 0.476 0.578
B.2 only seasonal alignment 0.127 0.212 0.367 0.432
B.3 only residual alignment 0.171 0.229 0.433 0.506
@@
Multi-level text alignment variants. Our results in Table 4 show that removing component alignment or prefixed prompts negatively impacts knowledge transfer during LLM reprogramming for effective time series forecasting. Specifically, without alignment (A.1), we observe a significant average performance drop of 75.4% across standard and few-shot forecasting tasks. We also examine the effect of aligning only two components to assess whether aligning just one component is sufficient in our multi-level alignment strategy. Retaining only seasonal alignment (B.2) achieves the best performance, though it still results in an average MSE increase of 4.5% across all scenarios. In contrast, keeping only trend alignment significantly degrades performance, with over 32.2% performance loss in both standard and few-shot tasks. Furthermore, altering the selection of anchors for trend alignment (C.1, C.2) increases MSE by over 14.5% when using noise anchors. Expanding the anchor selection with synonymous words produces results comparable to the default, with less than a 2% variation in MSE. Finally, removing component-specific instruction (D.1) and domain features (D.2) results in MSE increases of 7.7% and 1.7%, respectively.
Multi-level text alignment interpretation. We present a case study on ETTm1 using non-overlapping patching, where the patch stride is the same as the patch length, with different selected anchors shown in Figure 3. In our main experiments, the original anchors are “increase decrease upward downward linear exponential drift stable volatile stationary persistent rapid”. The attention map illustrates the optimized attention scores between input trend patches and aligned anchors. These matched anchors serve as textual shapelets for the time series tokens. Specifically, subplot (a) displays the optimized attention scores for synonymous anchors, which are consistent. The highlighted anchors, “rise”, “increase”, “climb”, “grow”, and “expand”, are associated with upward trend patches. In contrast, subplot (b) shows no highlighted anchor when all trend patches are aligned with noise words unrelated to time series trends. This case study demonstrates that aligned anchors effectively summarize the textual shapelets of the input trend patches.
@@ 5 Conclusion and future work
Appendix A Appendix
Our full results in few-shot forecasting tasks are detailed in Table 5. With the scope of 10% few-shot learning, our model’s secure SOTA performance in 13 out of 35 cases, spanning seven different time series benchmarks. Moreover, our model only lose to LLM4TS, which is neither interpretable nor light-weight.
Table 5: Full few-shot results on 10% training data.
Methods Ours Time-LLM LLM4TS GPT4TS DLinear PatchTST
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.629 0.548 0.530 0.492 0.417 0.432 0.458 0.456 0.492 0.495 0.516 0.485
192 0.720 0.534 0.671 0.546 0.469 0.468 0.570 0.516 0.565 0.538 0.598 0.524
336 0.893 0.622 0.907 0.639 0.505 0.499 0.608 0.535 0.721 0.622 0.657 0.550
720 1.210 0.772 0.917 0.647 0.708 0.572 0.725 0.591 0.986 0.743 0.762 0.610
Avg. 0.865 0.619 0.756 0.581 0.525 0.493 0.590 0.525 0.691 0.600 0.633 0.542
ETTh2 96 0.330 0.383 0.349 0.418 0.282 0.351 0.331 0.374 0.357 0.411 0.353 0.389
192 0.357 0.408 0.406 0.428 0.364 0.400 0.402 0.411 0.569 0.519 0.403 0.414
336 0.406 0.439 0.488 0.489 0.374 0.416 0.406 0.433 0.671 0.572 0.426 0.441
720 0.498 0.496 0.540 0.520 0.445 0.461 0.449 0.464 0.824 0.648 0.477 0.480
Avg. 0.397 0.431 0.446 0.464 0.366 0.407 0.397 0.421 0.605 0.538 0.415 0.431
ETTm1 96 0.360 0.389 0.297 0.349 0.360 0.388 0.390 0.404 0.352 0.392 0.410 0.419
192 0.429 0.431 0.336 0.373 0.386 0.401 0.429 0.423 0.382 0.412 0.437 0.434
336 0.446 0.465 0.362 0.390 0.415 0.417 0.469 0.439 0.419 0.434 0.476 0.454
720 0.489 0.495 0.410 0.421 0.470 0.445 0.569 0.498 0.490 0.477 0.681 0.556
Avg. 0.431 0.445 0.351 0.383 0.402 0.457 0.464 0.441 0.411 0.429 0.501 0.466
ETTm2 96 0.126 0.231 0.192 0.276 0.184 0.265 0.188 0.269 0.213 0.303 0.191 0.274
192 0.223 0.300 0.266 0.320 0.240 0.301 0.251 0.309 0.278 0.345 0.252 0.317
336 0.290 0.345 0.317 0.356 0.294 0.337 0.307 0.346 0.338 0.385 0.306 0.353
720 0.412 0.420 0.418 0.420 0.386 0.393 0.426 0.417 0.436 0.440 0.433 0.427
Avg. 0.262 0.324 0.292 0.343 0.276 0.324 0.293 0.335 0.316 0.368 0.296 0.343
Weather 96 0.102 0.177 0.164 0.220 0.158 0.207 0.163 0.215 0.171 0.224 0.165 0.215
192 0.164 0.234 0.215 0.258 0.204 0.249 0.210 0.254 0.215 0.263 0.210 0.257
336 0.230 0.281 0.259 0.294 0.254 0.288 0.256 0.292 0.258 0.299 0.259 0.297
720 0.334 0.363 0.319 0.326 0.322 0.336 0.321 0.339 0.320 0.346 0.332 0.346
Avg. 0.207 0.263 0.359 0.275 0.235 0.270 0.238 0.275 0.241 0.283 0.242 0.279
ECL 96 0.144 0.250 0.145 0.246 0.135 0.231 0.139 0.237 0.150 0.253 0.140 0.238
192 0.167 0.271 0.160 0.259 0.152 0.246 0.156 0.252 0.164 0.264 0.160 0.255
336 0.194 0.294 0.182 0.278 0.173 0.267 0.175 0.270 0.181 0.282 0.180 0.276
720 0.255 0.340 0.239 0.324 0.229 0.312 0.233 0.317 0.223 0.321 0.241 0.323
Avg. 0.190 0.288 0.182 0.277 0.172 0.264 0.176 0.269 0.180 0.280 0.180 0.273
Traffic 96 0.347 0.292 0.416 0.295 0.402 0.288 0.414 0.297 0.419 0.298 0.403 0.289
192 0.398 0.305 0.424 0.306 0.416 0.294 0.426 0.301 0.434 0.305 0.415 0.296
336 0.427 0.313 0.435 0.314 0.429 0.302 0.434 0.303 0.449 0.313 0.426 0.304
720 0.466 0.330 0.476 0.331 0.480 0.326 0.487 0.337 0.484 0.336 0.474 0.331
Avg. 0.409 0.310 0.438 0.312 0.432 0.303 0.440 0.310 0.447 0.313 0.430 0.305
[…]
Table 6: Full zero-shot learning results on ETT datasets.
Methods Ours Time-LLM GPT4TS LLMTime PatchTST DLinear
Datasets $\setminus$ Horizon MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 $\rightarrow$ ETTh2 96 0.263 0.337 0.264 0.340 0.335 0.374 0.510 0.576 0.304 0.350 0.347 0.400
192 0.315 0.374 0.332 0.376 0.412 0.417 0.523 0.586 0.386 0.400 0.447 0.460
336 0.371 0.414 0.395 0.424 0.441 0.444 0.640 0.637 0.414 0.428 0.515 0.505
720 0.435 0.462 0.423 0.459 0.438 0.452 2.296 1.034 0.419 0.443 0.665 0.589
Avg. 0.346 0.396 0.354 0.400 0.406 0.422 0.992 0.708 0.380 0.405 0.493 0.488
ETTh1 $\rightarrow$ ETTm2 96 0.191 0.296 0.224 0.311 0.236 0.315 0.646 0.563 0.215 0.304 0.255 0.357
192 0.259 0.338 0.270 0.339 0.287 0.342 0.934 0.654 0.275 0.339 0.338 0.413
336 0.317 0.370 0.336 0.378 0.341 0.374 1.157 0.728 0.334 0.373 0.425 0.465
720 0.409 0.424 0.410 0.422 0.435 0.422 4.730 1.531 0.431 0.424 0.640 0.573
Avg. 0.294 0.357 0.310 0.363 0.325 0.363 1.867 0.869 0.314 0.360 0.415 0.452
ETTh2 $\rightarrow$ ETTh1 96 0.585 0.510 0.541 0.503 0.732 0.577 1.130 0.777 0.485 0.465 0.689 0.555
192 0.677 0.554 0.559 0.515 0.758 0.559 1.242 0.820 0.565 0.509 0.707 0.568
336 0.700 0.562 0.620 0.551 0.759 0.578 1.328 0.864 0.581 0.515 0.710 0.577
720 0.693 0.579 0.729 0.627 0.781 0.597 4.145 1.461 0.628 0.561 0.704 0.596
Avg. 0.663 0.551 0.612 0.549 0.757 0.578 1.961 0.981 0.565 0.513 0.703 0.574
ETTh2 $\rightarrow$ ETTm2 96 0.181 0.288 0.218 0.304 0.253 0.329 0.646 0.563 0.226 0.309 0.240 0.336
192 0.235 0.324 0.265 0.335 0.293 0.346 0.934 0.654 0.289 0.345 0.295 0.369
336 0.294 0.357 0.327 0.370 0.347 0.376 1.157 0.728 0.348 0.379 0.345 0.397
720 0.395 0.411 0.401 0.416 0.446 0.429 4.730 1.531 0.439 0.427 0.432 0.442
Avg. 0.276 0.345 0.303 0.356 0.335 0.370 1.867 0.869 0.325 0.365 0.328 0.386
ETTm1 $\rightarrow$ ETTh2 96 0.415 0.437 0.331 0.383 0.353 0.392 0.510 0.576 0.354 0.385 0.365 0.415
192 0.486 0.477 0.353 0.399 0.443 0.437 0.523 0.586 0.447 0.434 0.454 0.462
336 0.397 0.433 0.400 0.428 0.469 0.461 0.640 0.637 0.481 0.463 0.496 0.494
720 0.451 0.475 0.417 0.448 0.466 0.468 2.296 1.034 0.474 0.471 0.541 0.529
Avg. 0.437 0.455 0.375 0.415 0.433 0.439 0.992 0.708 0.439 0.438 0.464 0.475
ETTm1 $\rightarrow$ ETTm2 96 0.081 0.185 0.194 0.270 0.217 0.294 0.646 0.563 0.195 0.271 0.221 0.314
192 0.162 0.250 0.243 0.304 0.277 0.327 0.934 0.654 0.258 0.311 0.286 0.359
336 0.250 0.311 0.295 0.341 0.331 0.360 1.157 0.728 0.317 0.348 0.357 0.406
720 0.376 0.392 0.367 0.385 0.429 0.413 4.730 1.531 0.416 0.404 0.476 0.476
Avg. 0.217 0.284 0.275 0.325 0.313 0.348 1.867 0.869 0.296 0.334 0.335 0.389
ETTm2 $\rightarrow$ ETTh2 96 0.485 0.473 0.322 0.369 0.360 0.401 0.510 0.576 0.327 0.367 0.333 0.391
192 0.496 0.481 0.359 0.396 0.434 0.437 0.523 0.586 0.411 0.418 0.441 0.456
336 0.542 0.509 0.439 0.452 0.460 0.459 0.640 0.637 0.439 0.447 0.505 0.503
720 0.437 0.463 0.448 0.468 0.485 0.477 2.296 1.034 0.459 0.470 0.543 0.534
Avg. 0.490 0.481 0.392 0.421 0.435 0.443 0.992 0.708 0.409 0.425 0.455 0.471
ETTm2 $\rightarrow$ ETTm1 96 0.345 0.377 0.446 0.415 0.747 0.558 1.179 0.781 0.491 0.437 0.570 0.490
192 0.518 0.462 0.496 0.452 0.781 0.560 1.327 0.846 0.530 0.470 0.590 0.506
336 0.735 0.541 0.507 0.463 0.778 0.578 1.478 0.902 0.565 0.497 0.706 0.567
720 0.653 0.533 0.556 0.482 0.769 0.573 3.749 1.408 0.686 0.565 0.731 0.584
Avg. 0.562 0.478 0.501 0.453 0.769 0.567 1.933 0.984 0.568 0.492 0.649 0.537

### v2

@@ 4 Experiments
ETTh1 0.424 0.448 0.420 0.438 0.404 0.418 0.428 0.426 0.422 0.437 0.413 0.435
ETTh2 0.378 0.408 0.364 0.403 0.333 0.383 0.355 0.394 0.431 0.446 0.330 0.379
ETTm1 0.251 0.325 0.351 0.383 0.343 0.378 0.352 0.383 0.357 0.378 0.351 0.380
ETTm2 0.214 0.289 0.270 0.328 0.251 0.313 0.267 0.326 0.267 0.333 0.255 0.315
Weather 0.171 0.226 0.229 0.267 0.223 0.260 0.237 0.271 0.248 0.300 0.225 0.264
ECL 0.159 0.262 0.164 0.261 0.159 0.253 0.167 0.263 0.166 0.263 0.161 0.252
Traffic 0.346 0.265 0.408 0.290 0.401 0.273 0.414 0.294 0.433 0.295 0.390 0.263
Zero-shot forecasting. Beyond few-shot learning, LLMs also show promise as effective zero-shot learners. In this section, we evaluate the zero-shot learning capabilities of the multi-level text-aligned LLM. Specifically, we assess how well the model performs on one dataset after being optimized on another. Similar to the few-shot learning setup, we use the long-term forecasting protocol and evaluate various cross-domain scenarios utilizing the ETT datasets.
Table 2: Zero-shot learning results on ETT datasets. All results are averaged from four different forecasting horizons: $H\in\{96,192,336,720\}$.
@@
The brief results are presented in Table 2. Our model demonstrates performance that is comparable to or surpasses other baselines. In data-scarce scenarios, our model significantly outperforms other LLM-based models, consistently providing better forecasts. Both our model and TimeLLM [6] outperform traditional baselines, likely due to cross-modality alignment, which more effectively activates LLMs’ knowledge transfer and reasoning capabilities for time series tasks. Additionally, our multi-level aligned embeddings better align language cues with temporal components of time series, enabling superior zero-shot forecasting performance compared to TimeLLM. More detailed experimental results for zero-shot forecasting are in the appendix and full paper [1].
Variant Long-term Few-shot
96 192 96 192
@@
B.1 only trend align 0.184 0.283 0.476 0.578
B.2 only seasonal align 0.127 0.212 0.367 0.432
B.3 only residual align 0.171 0.229 0.433 0.506
@@
Table 3: Comparison of different variants for long-term and few-shot forecasting on ETTm1 dataset with horizons 96 and 192.
Multi-level text alignment variants. Table 3 shows that removing component alignment or prefixed prompts hinders knowledge transfer in LLM reprogramming for time series forecasting. Without alignment (A.1), performance drops by 75.4% on average. Retaining only seasonal alignment (B.2) performs best but still increases MSE by 4.5%, while trend-only alignment causes a 32.2% loss. Using noise anchors (C.1) raises MSE by 14.5%, whereas synonymous anchors (C.2) yield similar results to the default with under 2% variation. Removing component-specific instruction (D.1) and domain features (D.2) increases MSE by 7.7% and 1.7%, respectively.
Multi-level text alignment interpretation. We conduct a case study on ETTm1 using non-overlapping patching, where the patch stride equals the patch length. Figure 3 shows the selected anchors, including “increase, decrease, upward, downward, linear, exponential, drift, stable, volatile, stationary, persistent, and rapid”. The attention map highlights optimized attention scores between trend patches and aligned anchors, serving as textual shapelets for time series tokens.
The attention map illustrates the optimized attention scores between input trend patches and aligned anchors. These matched anchors serve as textual shapelets for the time series tokens. Specifically, subplot (a) displays the optimized attention scores for synonymous anchors, which are consistent. The highlighted anchors, “rise”, “increase”, “climb”, “grow”, and “expand”, are associated with upward trend patches. In contrast, subplot (b) shows no highlighted anchor when all trend patches are aligned with noise words unrelated to time series trends. This case study demonstrates that aligned anchors effectively summarize the textual shapelets of the input trend patches.
@@ 5 Conclusion and future work
6 Acknowledgments
This work was supported in part by the NSF under Grant No. 1943486, No. 2246757, No. 2315612, and 1946231.